end of the current year, ARS Industries has the following account balances before making an adjusting entry for uncollectible accounts Accounts Receivable has a balance $1,300,000; Allowance for Doubtful Accounts $ 7,000 Credit sales for this year $8,500,000 A. If the company uses the percent of sales method and estimates that ½ of 1% of credit sales for the year will be uncollectible, then a. Determine the bad debt expense for the period, and journalize the adjusting entries. (2 points) b. Determine the adjusted balances of Accounts Receivable, Allowance for Doubtful Accounts, and Net Realizable Value of Accounts Receivable. (3 points) Adjusted Balance Accounts Receivable _____________ Allowance for Doubtful Accounts _____________ Net Realizable Value of Accounts Receivable _____________ B. An aging of the company’s accounts receivable on December 31, and a historical analysis of the percentage of uncollectible accounts in each age class follow: Age Class Balance Percent Uncollectible Not past due $715,000 1% 1-30 days 310,000 2% 31-60 days 102,000 15% 61-90 days 76,000 30% 91-120 days 97,000 60% $1,300,000 a. Determine the bad debt expense for the period, and journalize the adjusting entries. (2 points) b. Determine the adjusted balances of Accounts Receivable, Allowance for Doubtful Accounts, and Net Realizable Value of Accounts Receivable. (3 points) Adjusted Balance Accounts Receivable _____________ Allowance for Doubtful Accounts _____________ Net Realizable Value of Accounts Receivable ____________
Answer #1

1a) Bad debt expense = 8500000*1%*1/2 = 42500

Adjusting entry

Account title and explanation Debit Credit
bad debt expense 42500
Allowance for doubtful accounts 42500

1b) Adjusted balance

Account receivable 1300000
Less: Allowance for doubtful account -49500
Net account receivable, net 1250500

2a) Schedule

Days Account receivable % Uncollectible $ Uncollectible
Not past due 715000 1% 7150
1-30 days 310000 2% 6200
31-60 days 102000 15% 15300
61-90 days 76000 30% 22800
91-120 days 97000 60% 58200
Total 109650

Bad debt expense = 109650-7000 = 102650

Account title and explanation Debit Credit
bad debt expense 102650
Allowance for doubtful accounts 102650

2b) Adjusted balance

Account receivable 1300000
Less: Allowance for doubtful account -109650
Net account receivable, net 1190350